# SHINE: Social Homology Identification for Navigation in Crowded Environments

### v1

@@ 5 Navigation system
In the following, we present a navigation framework which derives topologically distinct trajectories, selects one guidance trajectory using the learned cost function, and applies a local motion planner to track it. A schematic representation of the three main components of the system is provided in Figure 4.
Figure 4: The guidance proposal algorithm of Section 5.1 finds guidance trajectories that are topologically distinct. The guidance selection component of Section 4 selects the one that defines the high-level behavior of the robot, and an MPC-based planner Section 5.2 computes the motion commands based on the selected guidance.
@@ 5.2 Local Optimization-based Planner
The selected guidance trajectory identifies the most suitable homology class but is not guaranteed to be dynamically feasible or collision free. To obtain a high-quality trajectory, we initialize and track the guidance trajectory with a local optimization-based planner that refines the trajectory in the same state space, similar to (de Groot et al., 2023). The local planner is an MPC that solves a trajectory optimization, enforcing dynamic and collision avoidance constraints and returns a smooth trajectory near the guidance trajectory. We use the planner LMPCC (Brito et al., 2019). From the guidance trajectory, we first derive a guidance path, $[0,S]\to\mathbb{R}^{2}$ that maps the distance along the path (of length $S$) to an $x,y$-position. The MPC solves the optimization problem
@@ 6.1 Experimental setup
Hardware setup. The real-world evaluation is conducted on a Clearpath Jackal mobile robot in a room of size 7.5 m x 6.5 m. The platform is equipped with an Intel i5 CPU@2.6GHz. The robot and pedestrian positions are determined through a marker-based tracking system. Pedestrian predictions are based on the assumption of constant velocity, with the velocity obtained using a Kalman filter. Note that we decouple perception from navigation, and in the simulation experiments this constant velocity assumption is not applied.
@@ 6.2 Prediction evaluation
  • Minimum acceleration: The smoothest trajectory has the minimum cost: $\mathcal{J}_{acc}=\sum_{i\in N_{\tau}}\alpha^{i}||\bm{a}_{i}||$, where $\alpha\approx 1$ to discount accelerations in time.
@@
$\mathcal{H}_{right}(\bm{\tau})=[h_{0}(\bm{\tau}_{right}\sqcup-\bm{\tau}),\dots,h_{M}(\bm{\tau}_{right}\sqcup-\bm{\tau})]^{T},$ (17)
@@
The accuracy results are shown in Table 4. Our method significantly outperforms the other approaches. We noticed that, once the human makes a decision on how to pass other humans, it usually sticks with it, following a previously made decision. This explains why all methods achieve good accuracy ($>0.5$). The results thus show that our method does not only select just the homology class the human was previously following, but also accurately estimates the decision at any time step.
@@ 6.3 Simulation experiments
In these scenarios, both LMPCC-based planners (LMPCC and SHINE) are safer (i.e. higher success rates and lower collision rates). Social Force’s goal is navigating in a way that respects social norms with physics-inspired principles of attractive and repulsive forces, and its simplicity compared to the other planners could contribute to a higher collision rate. The DRL planner was trained in CrowdNav simulator, which uses random positions for humans. Its success rate could increase after training it in corridor scenarios, but it would mean that its effectiveness does not translate to different settings. The metrics of average path length, time to reach the goal, speed, distance to obstacles and time to collision are similar and comparable for all the planners; as their meaning is more related to the local constraints, and we set their parameters to behave similarly.
@@ 6.4 Real world experiments
We evaluated our approach in the real platform and compared it with other methods of the state of the art. We conducted experiments to see the limitations of other methods and the behavior of ours in those situations. Particularly, we observed problems of Social Force in what we called push out scenarios, Social DRL in scenarios different from what experienced in training and LMPCC in zig-zag scenarios. In addition, we conducted experiments to test the navigation capabilities in crowded environments. In every experiment, the robot is placed in a corner of the room and has to navigate to the other corner. A visual summary of the real-world experiments is recorded in the supplementary video.
@@ 6.4.4 Crowded scenarios
We conducted experiments with five people randomly walking around the room with the robot navigating with Social Force, the DRL planner, the Guidance-MPCC, and SHINE; for five minutes each. We told people to walk normally, assuming the robot was going to behave like another person. People did not know what algorithm was working at every moment (they were randomized). We asked the participants to rate how conformable they had been after interacting with each planner, and all of them assigned higher grades to SHINE than to Guidance-MPCC. Social Forces and DRL had a very reactive behavior that was not desirable for navigating around humans, even leading to some collisions.
@@ 7 Discussion
The results of this work hold remarkable implications for the field of social navigation. One of the key implications is the ability to learn high-level discrete decisions from humans, unlike most of the learning-based social navigation approaches, which focus on imitating the low level behavior of humans (like DRL or imitation learning). They do not have collision constraints, and their performance is limited to what they have learned. Another problem that some learning-based approaches have is that they need a simulator to train the model, so their performance is limited to the realism of the simulations. With SHINE, however, in the worst case, the network will select a collision-free path that goes through a non-social homology class, still leading to a smooth and collision-free trajectory. Moreover, our work could be easily integrated with local optimization-based planners that include social costs in the optimization, achieving local social motion too. Therefore, our method is safer and easier to train compared to pure learning-based approaches, while we still incorporate learned social aspects to the navigation problem.
@@
SHINE also provides a combination of global and local planning in dynamic environments. Unlike only reactive and local planners, it may change between homology classes in real time. This is specially important in crowded environments, as human decisions may change in time and the robot behavior should adapt to those changes.
@@ 8 Conclusion
The author(s) disclosed receipt of the following financial support for the research, authorship, and/or publication of this article: This work was supported by the Ministry of Science, Innovation and Universities of Spain through the research project MCIN/AEI/10.13039/501100011033/FEDER-UE, from the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 101017008, and from the European Union (ERC, INTERACT, 101041863). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.

### v3

@@ 5 Navigation system
In the following, we present a navigation framework which derives topologically distinct trajectories, selects one guidance trajectory using the learned cost function, and applies a local motion planner to track it. A schematic representation of the three main components of the system is provided in Fig. 5.
Figure 5: Schematic representation of SHINE. First, the guidance proposal algorithm in Section 5.1 finds trajectories (two blue lines originating from the robot) in different homology classes ($\mathcal{T}^{*}_{k}$). Second, each trajectory is compared to a reference trajectory to identify its homology class. Third, the guidance selection component in Section 4, that was trained to select socially compliant trajectories from human motion data, selects the guidance trajectory ($\tau_{k}$). Finally, the MPC-based planner in Section 5.2 computes the motion commands based on $\tau_{k}$.
@@ 5.2 Local Optimization-based Planner
The selected guidance trajectory identifies the most suitable homology class but is not guaranteed to be dynamically feasible or collision free. To obtain a high-quality trajectory, we initialize and track the guidance trajectory with a local optimization-based planner that refines the trajectory in the same state space, similar to (de Groot et al., 2023). The local planner is an MPC that solves a trajectory optimization problem, enforcing dynamic and collision avoidance constraints and returns a smooth trajectory near the guidance trajectory. We use the planner LMPCC (Brito et al., 2019). From the guidance trajectory, we first derive a guidance path, $[0,S]\to\mathbb{R}^{2}$ that maps the distance along the path (of length $S$) to an $x,y$-position. The MPC solves the optimization problem
@@ 6.1 Experimental setup
Hardware setup. The real-world evaluation is conducted on a Clearpath Jackal mobile robot in a room of size 7.5 m x 6.5 m. The platform is equipped with an Intel i5 CPU@2.6GHz. The robot and pedestrian positions are determined through a marker-based tracking system. Pedestrian predictions are based on the assumption of constant velocity, with the velocity obtained using a Kalman filter.
@@ 6.2 Prediction evaluation
  • Minimum acceleration: The smoothest trajectory has the minimum cost: $\mathcal{J}_{acc}=\sum_{i\in N_{\tau}}\alpha^{i}||\bm{a}_{i}||$, where $\alpha\approx 1$ to discount accelerations in time (0.95 in the implementation).
@@
$\mathcal{H}_{right}(\bm{\tau})=[h_{1}(\bm{\tau}_{right}\sqcup-\bm{\tau}),\dots,h_{M}(\bm{\tau}_{right}\sqcup-\bm{\tau})]^{T},$ (16)
@@
The accuracy results are shown in Table 4. Our method significantly outperforms the other approaches. We noticed that, once the human makes a decision on how to pass other humans, it usually sticks with it, following a previously made decision. This explains why all methods achieve good accuracy (>0.5>0.5> 0.5). The results thus show that our method does not only select just the homology class the human was previously following, but also accurately estimates the decision at any time step.
@@ 6.3 Simulation experiments
In these scenarios, both LMPCC-based planners (LMPCC and SHINE) are safer with higher success rates (number of times the robot completes the task without collisions or getting stuck per scenario) and lower collision rates (number of times the robot collides per scenario). Social Force’s goal is navigating in a way that respects social norms with physics-inspired principles of attractive and repulsive forces, and its simplicity compared to the other planners could contribute to a higher collision rate. The DRL planner was trained in CrowdNav simulator, which uses random positions for humans. Its success rate could increase after training it in corridor scenarios, but it would mean that its effectiveness does not translate to different settings. The metrics of average path length, time to reach the goal, speed, distance to obstacles and time to collision are similar and comparable for all the planners; as their meaning is more related to the local constraints, and we set their parameters to behave similarly.
@@
Finally, we conducted experiments to assess the benefit of learning the guidance trajectories from humans rather than designing a hand-crafted cost for it. To do so, we used SocNavBench (Biswas et al., 2022), a simulator that generates humans using prerecorded trajectories, to compare SHINE with Guidance-MPCC using $\mathcal{J}_{len}$, $\mathcal{J}_{acc}$ and $\mathcal{J}_{mix}$. In this way, we intend to have more realistic simulations that resemble the real world and to use an existing benchmark. The metrics gathered for 5 runs of the 33 scenarios proposed as a benchmark in SocNavBench are shown in Table 7. We do not include comparisons with other planners of the simulator, as they do not consider robot dynamics. In this simulator, humans are spawned at the time when the recorded data of each specific human starts. Therefore, humans are sometimes spawned right into the robot position, resulting in an unavoidable collision and not reliable metrics. In addition, some scenarios may not have trajectories in different homology classes as either the robot faces no obstacles in its way to the goal or the scenario is too crowded. Therefore, differences in Table 7 among the methods are small. Thus, we selected the benchmark scenarios where SHINE and $\mathcal{J}_{acc}$, which has the best performance in Table 7 among the hand-crafted costs, navigated within different homology classes without collisions. The resulting metrics in those eleven scenarios are shown in Table 8. The results prove a better performance of SHINE with respect any other alternative used as baseline when it chooses a different homology class. One example is shown in Fig. 10, where both G-MPCC and SHINE reach the goal succesfully at similar times, but SHINE in a way that is less intrusive regarding pedestrians. The metrics show that SHINE exhibits safety with high success rates, efficiency with low task duration and path length ratio metrics, as well as higher minimum distance to pedestrians and minimum time to collision metrics per episode. It is specially remarkable how it reaches the goal in considerably less time while keeping the highest time to collision to pedestrians.
Table 7: Metrics gathered in SocNavBench.
Metric $\mathcal{J}_{len}$ $\mathcal{J}_{acc}$ $\mathcal{J}_{mix}$ SHINE
Success rate $\uparrow$ 0.850 0.912 0.856 0.912
Duration (s) $\downarrow$ 12.676 (2.7) 12.211 (2.5) 12.663 (2.7) 11.994 (2.6)
Path lenght ratio $\downarrow$ 0.976 (0.0) 0.964 (0.0) 0.976 (0.0) 0.957 (0.0)
Min. closest ped. dist. (m) $\uparrow$ 0.880 (0.9) 0.942 (0.8) 0.879 (0.8) 0.945 (0.8)
Min. time to coll. $\uparrow$ 1.065 (2.1) 1.114 (2.0) 1.039 (2.1) 1.141 (2.0)
Table 8: Metrics gathered in SocNavBench in scenarios where SHINE and Guidance-MPCC with $\mathcal{J}_{acc}$ are in different homology classes.
Metric $\mathcal{J}_{len}$ $\mathcal{J}_{acc}$ $\mathcal{J}_{mix}$ SHINE
Success rate $\uparrow$ 0.909 1.000 0.909 1.000
Duration (s) $\downarrow$ 13.735 (3.4) 13.223 (2.1) 13.740 (2.7) 12.527 (2.6)
Path lenght ratio $\downarrow$ 0.993 (0.0) 0.992 (0.0) 1.016 (0.1) 0.972 (0.0)
Min. closest ped. dist. (m) $\uparrow$ 0.523 (0.0) 0.522 (0.0) 0.517 (0.0) 0.523 (0.0)
Min. time to coll. $\uparrow$ 0.401 (0.2) 0.368 (0.2) 0.360 (0.2) 0.444 (0.3)
Figure 10: A scenario (named univ_trapTL) of SocNavBench showing SHINE and G-MPCC with $\mathcal{J}_{acc}$ cost trajectories.
@@ 6.4 Real world experiments
We evaluated our approach in the real platform and compared it with other methods of the state of the art. We conducted two kind of experiments. First, a group of experiments to assess the limitations of other methods and the behavior of SHINE in those situations. We designed simple scenarios which could make planners fail due to their design. Particularly, we observed problems of Social Force in what we called push out scenarios, Social DRL in scenarios different from what experienced in training and LMPCC in zig-zag scenarios. The second type of experiments were conducted to test the navigation capabilities in crowded environments. In every experiment, the robot is placed in a corner of the room and has to navigate to the other corner. A visual summary of the real-world experiments is recorded in the supplementary video, which may be accessed in the project repository (Martinez-Baselga et al., 2024).
@@ 6.4.4 Crowded scenarios
This experiment differed from the others in that we were not addressing a specific limitation of a method but rather assessing overall performance. We conducted the experiment in scenarios involving several moving pedestrians. Five participants were randomly selected and instructed to wander around the room freely. They were informed that a robot would also be moving in the room and that they should assume it would behave like another person.
To begin, participants were asked to walk around the room for one minute without the robot, allowing them to get accustomed to the environment. Following this, they repeated the exercise while sharing the space with the moving robot for five minutes. Afterward, each participant individually and separately evaluated the robot’s motion qualitatively for one minute to avoid mutual influence.
This sequence was repeated four times, each time with a different motion planner controlling the robot. The planners tested in the experiment were Social Force, DRL, Guidance-MPCC, and SHINE, in that order. We recorded the positions of both the robot and the pedestrians, participants’ evaluations of each planner, and a video of the experiment.
In the evaluations, all the participants assigned higher grades to SHINE than to Guidance-MPCC. Social Force and DRL had a very reactive behavior that was not desirable for navigating around humans, even leading to some collisions.
@@ 7 Discussion
The results of this work hold remarkable implications for the field of social navigation. One of the key implications is the ability to learn high-level discrete decisions from humans, unlike most of the learning-based social navigation approaches, which focus on imitating the low level behavior of humans (like DRL or imitation learning). They do not have collision constraints, and their performance is limited to the scenarios and behaviors seen during training, which may not exhibit all the intricacies of the real world, being extremely sensitive to out-of-distribution data. Another problem that some learning-based approaches have is that they need a simulator to train the model, so their performance is limited to the realism of the simulations. With SHINE, however, in the worst case, the network will select a collision-free path that goes through a non-social homology class, still leading to a smooth and collision-free trajectory. Moreover, our work could be easily integrated with local optimization-based planners that include social costs in the optimization, achieving local social motion too. Therefore, our method is safer and easier to train compared to pure learning-based approaches, while we still incorporate learned social aspects to the navigation problem.
@@
SHINE provides a combination of global and local planning in dynamic environments. Unlike only reactive and local planners, it may change between homology classes in real time. This is specially important in crowded environments, as human decisions may change in time and the robot behavior should adapt to those changes.
The homology class selection module relies solely on the previous positions of obstacles and the robot and has been trained using real-world human trajectory data, implicitly accounting for possible interactions with humans and among humans themselves. Through this approach, SHINE implicitly couples prediction and planning, using the learned knowledge in the planning process. Although the local planner, LMPCC, does not integrate prediction and planning, it generates a local plan within that homology class, thereby accounting for human interactions at a global level.
The prediction module does not try to find an optimal homology class, as it would depend on the user-specified navigation objective of the robot and it may even not exist. The goal of the module is finding the homology class that is the most likely to be chosen by a human. Due to the natural multi-modality in human trajectory selection, achieving a perfect prediction accuracy is impossible, having scenarios where more than one homology class are equally good depending on specific human preferences. Nevertheless, this is not a problem for SHINE, as selecting any equally good homology class results in equally good high-level motion planning.
@@ 8 Conclusion
Even though SHINE accounts for human future interactions in a high level, the local planner used, LMPCC, does not plan considering them. Future work could involve studying how to integrate a local planner that couples low level interactions with local planning. In addition, extending the training data with datasets that involve other scenarios such as groups of people could expand and improve SHINE performance and adaptability with more kinds of social interactions.
@@
The author(s) disclosed receipt of the following financial support for the research, authorship, and/or publication of this article: This work was supported by MICIU/AEI/10.13039/501100011033 and ERDF/EU under grants (PID2022-139615OB-I00 and PRE2020-094415), from the Government of Aragón under grant DGA T45-23R, from the European Union’s Horizon 2020 research and innovation programme under grant agreement No. (101017008), and from the European Union ERC, INTERACT, (101041863). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.